# Lattice points on determinant surfaces and the spectrum of the automorphic Laplacian

### v1

@@ 9.2. Analysis of $R_{3,w}(X)$
for a suitable $\varepsilon>0$. By Prop. 5.5, we have,
$\frac{e\left(\frac{1}{8}\right)}{\sqrt{2}}$ $\int\limits_{X}^{X+H}\frac{w(u)}{u^{1/4}}\mathop{\sum_{l}\sum_{m}\sum_{n}}_{nl\leq m\frac{u}{X}}\sum_{0<c\leq C_{0}^{\star}}\frac{S(nr_{1},-m,c)}{c}\left(\frac{c}{mn}\right)^{\frac{1}{4}}e\left(-2\sqrt{\frac{mnu}{c}}\right)w\left(\frac{mc}{n}\frac{u}{X}\right)\ll X^{\frac{3}{4}+\varepsilon}\left(\frac{X}{H}\right)^{\frac{3}{2}}.$ (9.7)
@@
$R_{3}(M,N,C,X,u)$ $:=\mathop{\sum_{m}\sum_{n}}_{nl\leq m\frac{u}{X}}b_{1}\left(\frac{m}{M}\right)b_{2}\left(\frac{n}{N}\right)\sum_{c}\frac{S(nr_{1},-m,c)}{c}e\left(-2\sqrt{\frac{mnu}{c}}\right)w\left(\frac{mc}{n}\frac{u}{X}\right)V\left(\frac{c}{C}\right)g^{\star}(m,n,c),$
@@
We then have the following proposition which can be compared with Prop. 9.1.
@@
We give the details of the proof of Prop. 9.1 below and the proof of Prop. 9.2 is almost identical and we only give a very brief sketch of it at the end.
@@ 10.1. Decomposition of $R_{1}(M,N,C,X)$
say, where
$\check{f}_{m,n,C}(\eta)=\frac{4}{\pi}\int\limits_{0}^{\infty}K_{2i\eta}(t)f_{m,n,C}(t)\frac{\textnormal{d}t}{t}.$
From the support of $V$, the argument $t$ of $f_{m,n,C}(t)$ (see (10.1)) is very small. Indeed, it lies in the the interval
@@
since $C>C_{0}$ (see (9.4)), $H\geq\sqrt{X}$ (see (3.3)) and $r=O(X^{1/3})$ (see Remark 16.1). We now expand $K_{2i\eta}(t)$ as a power series:
@@
$\mathcal{D}_{\pm}(M,N,C):=\mathop{\sum_{m}\sum_{n}}_{nl\leq m}b_{1}\left(\frac{m}{M}\right)b_{2}\left(\frac{n}{N}\right)\sum_{j=1}^{\infty}\rho_{j}(nr_{1})\rho_{j}(m)\check{f}_{m,n,C,\pm}(\kappa_{j}),$ (10.3)
$\mathcal{C}_{\pm}(M,N,C)$ $:=\frac{1}{\pi}\mathop{\sum_{m}\sum_{n}}_{nl\leq m}b_{1}\left(\frac{m}{M}\right)b_{2}\left(\frac{n}{N}\right)\int\limits_{-\infty}^{\infty}(nmr_{1})^{-i\eta}\sigma_{2i\eta}(nr_{1})\sigma_{2i\eta}(m)\frac{\cosh(\pi\eta)\check{f}_{m,n,C,\pm}(\eta)}{\left|\zeta(1+2i\eta)\right|^{2}}\textnormal{d}\eta,$ (10.4)
@@
$\check{f}_{m,n,C,\pm}(\eta)$ $:=\frac{4}{\pi}\int\limits_{0}^{\infty}K_{2i\eta,\pm}(t)f_{m,n,C}(t)\frac{\textnormal{d}t}{t}.$ (10.5)
@@
$I_{\pm}(m,n,C,\nu,\eta,r_{1})$ $:=\int t^{\mp 2i\eta+2\nu}w\left(\frac{4\pi m\sqrt{mnr_{1}}}{nt}\right)V\left(\frac{4\pi\sqrt{mnr_{1}}}{tC}\right)g\left(m,n,\frac{4\pi\sqrt{mnr_{1}}}{t}\right)e\left(\frac{-X^{\frac{1}{2}}t^{\frac{1}{2}}(mn)^{\frac{1}{4}}}{\sqrt{\pi}r_{1}^{1/4}}\right)\frac{\textnormal{d}t}{t}.$
@@ 11. Estimation of $\mathcal{D}_{+}(M,N,C)$ and $\mathcal{C}_{+}(M,N,C)$
Here, the phase function $-h_{+}$ has no stationary point and the first derivative bound together with the Weyl law suffices.
@@
$\check{f}_{m,n,C,+}(\eta)\ll\begin{cases}\left(\frac{C}{N^{2}l}\right)^{\frac{3}{4}}\frac{1}{\eta^{\frac{3}{2}}}e^{-\pi|\eta|}&\text{if }\eta>0,\\ X^{-100},&\text{if }\eta\gg\frac{MCX^{\varepsilon}}{NH}.\end{cases}$
@@ 11.1. Estimation of $\mathcal{D}_{+}(M,N,C)$
$\Theta_{c_{1},c_{2}}(M,N,K_{1},K_{2},\lambda)$ $:=\sum_{\begin{subarray}{c}c_{1}K_{1}\leq\kappa_{j}\leq c_{2}K_{2}\end{subarray}}\mathop{\sum_{m}\sum_{n}}b_{1}\left(\frac{m}{M}\right)b_{2}\left(\frac{n}{N}\right)|\rho_{j}(nq)\rho_{j}(m)|\frac{e^{-\pi|\kappa_{j}|}}{\kappa_{j}^{\lambda}}$
$\ll\begin{cases}q^{\theta+\varepsilon}(MN)^{1+\varepsilon}K_{2}^{2-\lambda+\varepsilon}&\text{if}\quad 0\leq\lambda\leq 2\\ q^{\theta+\varepsilon}(MN)^{1+\varepsilon}\text{if}\quad\lambda>2.\end{cases}$
@@
We subdivide the $j$-sum into dyadic pieces with $\kappa_{j}\asymp K$ such that $c_{1}K_{1}\leq K\leq c_{2}K_{2}$, and apply the Cauchy-Schwarz inequality to the $m$-sum and the $n$-sum, getting
@@
Since we know from Prop. 11.1 that the integral $I_{+}(m,n,C,\nu,\eta,r_{1})$ (see (10.8)) is negligibly small for $\kappa_{j}\gg\frac{MCX^{\varepsilon}}{NH}$, it is enough to estimate $\mathcal{D}_{+}(M,N,C)$ (see (10.3) and (10.7)) for $1\ll\kappa_{j}\ll\frac{MCX^{\varepsilon}}{NH}$. Using (10.3), Prop. 11.1, Prop. 11.3, and (9.5), we easily get
@@ 11.2. Estimation of $\mathcal{C}_{+}(M,N,C)$
$\Theta(M,N,K,\lambda)$ $:=\mathop{\sum_{m}\sum_{n}}_{nl\leq m}b_{1}\left(\frac{m}{M}\right)b_{2}\left(\frac{n}{N}\right)\int\limits_{0}^{\infty}(nr_{1}m)^{-i\eta}\sigma_{2i\eta}(nr_{1})\sigma_{2i\eta}(m)\frac{e^{-\pi|\eta|}\cosh(\pi\eta)}{\eta^{\lambda}\left|\zeta(1+2i\eta)\right|^{2}}\textnormal{d}\eta$
@@
We know from [Tit86, (2.1.16)] that for $0<\eta<\varepsilon$,
$\eta^{\lambda}\zeta(1+2i\eta)^{2}=\eta^{\lambda}\left(\frac{1}{2i\eta}+\gamma+O\left(\eta\right)\right)^{2}\gg\eta^{\lambda-2}.$
Also, from [Tit86, (3.6.5)], for $\eta\geq\varepsilon$,
$\frac{1}{\left|\zeta(1+2i\eta)\right|}\ll_{\varepsilon}|\log\eta|^{7}.$
@@
This is clear from Prop. 11.1, (10.4) and Prop. 11.5. ∎
@@ 12. Estimation of $\mathcal{D}_{-}(M,N,C)$ for $C_{0}\leq C\leq C_{1}$
Our next task is to estimate $\mathcal{D}_{-}(M,N,C)$ and $\mathcal{C}_{-}(M,N,C)$ (see (10.3), (10.4) and (10.6)). The sum $\mathcal{D}_{-}(M,N,C)$ (see (10.3)) is considerably more difficult to estimate than $\mathcal{D}_{+}(M,N,C)$ because the phase function $-h_{-}(u)$ (see (10.10)) can have a stationary point inside the range of the integral $I_{-}(m,n,C,\nu,\eta,r_{1})$ (see (10.8)). Differentiating $h_{-}(u)$ (see (10.10)),
@@ 12.1. Further truncation of the $c$-sum
From (3.5) and the fact that $V$ is supported in $[1,4]$, it follows that $w(\frac{4\pi m\sqrt{mn}}{nu^{2}})=1$ if $C_{0}<C\leq C_{1}$. In this range of $C$, the integral we need to analyze becomes
$I_{-}(m,n,C,\nu,\eta,r_{1})=2r_{1}^{\nu+i\eta}\int_{{\mathbb{R}}}A(u)e(-h_{-}(u))\textnormal{d}u,$
where
$A(u)=\frac{u^{4\nu}}{u}V\left(\frac{4\pi\sqrt{mn}}{u^{2}C}\right)g\left(m,n,\frac{4\pi\sqrt{mn}}{u^{2}}\right).$
@@
Unlike the previous case of $\mathcal{D}_{+}(M,N,C)$, an application of the Weyl law is not sufficient here. We need to obtain an explicit asymptotic expansion of the integral $I_{-}(m,n,C,\nu,\eta,r_{1})$ by applying Lemma 2.9. This is essential because if we estimate $I_{-}(m,n,C,\nu,\eta,r_{1})$ then our final bound will not be satisfactory. We will see that it is enough to consider the leading term $\mathcal{L}(m,n,\kappa_{j},\nu)$ of this asymptotic expansion (see (LABEL:lt). The phase $e(-h(u_{0}))$ occurring in $\mathcal{L}(m,n,\kappa_{j},\nu)$ simplifies to $(mn)^{-i\kappa_{j}}$, up to multiplication by a non-oscillatory function that does not involve $m$ and $n$, and we need to take advantage of the cancellations arising out of this oscillatory function. In order to implement this idea we will first have to separate the variables $m$ and $n$ which are at the moment entangled in the function $V$ and they are also constrained by the condition $nl\leq m$. Dropping the condition $nl\leq m$ by positivity as we had done earlier will be too costly here. We remove this condition by approximating the indicator function $\mathbf{1}_{[1,\infty)}(m/nl)$ by the Perron formula (Lemma 2.4). Furthermore, the variables $m$ and $n$ occurring in the argument of the function $V$ are separated by Mellin inversion. Finally, after the two variables are separated, we apply the Cauchy-Schwarz inequality and the the Large Sieve inequality Lemma 2.16 for twisted Fourier coefficients $\rho_{j}(n)n^{-i\kappa_{j}}$.
@@
where $A$ and $\delta$ are arbitrary subject to the conditions $A>0$ and $100<\delta<1/$ 10. We define
@@
$\mathcal{D}_{-,k}(M,N,C)$ $:=4\mathop{\sum_{m}\sum_{n}}_{nl\leq m}b_{1}\left(\frac{m}{M}\right)b_{2}\left(\frac{n}{N}\right)\sum_{j=1}^{\infty}\rho_{j}(nr_{1})\rho_{j}(m)\frac{1}{\sinh(2\pi\kappa_{j})\Gamma(1+2i\kappa_{j})}$ (12.8)
×$\times\sum_{\nu=0}^{\infty}\frac{2^{-2\nu}}{\nu!(1+2i\kappa_{j})\dots(\nu+2i\kappa_{j})}\left[\frac{2r_{1}^{\nu+i\eta}}{\sqrt{h^{\prime\prime}_{-}(u_{0})}}e\left(-h_{-}(u_{0})-\frac{1}{8}\right)p_{k}(u_{0})\right],$
@@
$\mathcal{D}_{-,Z}(M,N,C)$ $:=4\mathop{\sum_{m}\sum_{n}}_{nl\leq m}b_{1}\left(\frac{m}{M}\right)b_{2}\left(\frac{n}{N}\right)\sum_{j=1}^{\infty}\left|\frac{\rho_{j}(nr_{1})\rho_{j}(m)\kappa_{j}^{-A}}{\sinh(2\pi\kappa_{j})\Gamma(1+2i\kappa_{j})}\right|$ $\sum_{\nu=0}^{\infty}\frac{2^{-2\nu}r_{1}^{\nu}}{\nu!\left|(1+2i\kappa_{j})\dots(\nu+2i\kappa_{j})\right|}.$ (12.9)
@@
$\mathcal{L}(m,n,\kappa_{j},\nu,C):=2\frac{A(u_{0})}{\sqrt{h^{\prime\prime}_{-}(u_{0})}}e\left(-h_{-}(u_{0})-\frac{1}{8}\right)$ (12.10)
@@
$\frac{\pi}{2}\left(\frac{mnX}{C}\right)^{\frac{1}{2}}\leq\kappa_{j}\leq\pi\left(\frac{mnX}{C}\right)^{\frac{1}{2}};$ (12.11)
in particular, we have
$\frac{\pi}{2}N\left(\frac{Xl}{C}\right)^{\frac{1}{2}}\leq\kappa_{j}\leq 4\pi\left(\frac{MNX}{C}\right)^{\frac{1}{2}}.$ (12.12)
@@
$\mathcal{D}_{-,0}(M,N,C):=4\mathop{\sum_{m}\sum_{n}}_{nl\leq m}b_{1}\left(\frac{m}{M}\right)b_{2}\left(\frac{n}{N}\right)T(m,n,C),$ (12.13)
@@
$T(m,n,C):=\sum_{\nu=0}^{\infty}\sum_{\frac{\pi}{2}N\left(\frac{Xl}{C}\right)^{\frac{1}{2}}\leq\kappa_{j}\leq 4\pi\left(\frac{MNX}{C}\right)^{\frac{1}{2}}}\rho_{j}(nr_{1})\rho_{j}(m)\mathcal{V}(\kappa_{j},\nu,r_{1})\mathcal{L}(m,n,\kappa_{j},\nu,C),$
@@
$\mathcal{V}(\kappa_{j},\nu,r_{1})$ $:=\frac{1}{\sinh(2\pi\kappa_{j})}\frac{2^{-2i\kappa_{j}-2\nu}r_{1}^{\nu+i\kappa_{j}}}{\Gamma(1+2i\kappa_{j})\nu!(1+2i\kappa_{j})\dots(\nu+2i\kappa_{j})}\ll\frac{e^{-\pi\kappa_{j}}r_{1}^{\nu}}{\nu!\sqrt{\kappa_{j}}},$ (12.14)
@@
$\tilde{T}(m,n,C):=\sum_{\nu=0}^{\infty}\sum_{\frac{\pi}{2}N\left(\frac{Xl}{C}\right)^{\frac{1}{2}}\leq\kappa_{j}\leq 4\pi\left(\frac{MNX}{C}\right)^{\frac{1}{2}}}|\rho_{j}(nr_{1})\rho_{j}(m)\mathcal{V}(\kappa_{j},\nu,r_{1})\mathcal{L}(m,n,\kappa_{j},\nu,C)|.$ (12.15)

### v2

@@ 8.2. Analysis of $R_{3,w}(X)$
for a suitable $\varepsilon>0$. By Prop. 5.4, we have,
$\frac{e\left(\frac{1}{8}\right)}{\sqrt{2}}$ $\int\limits_{X}^{X+H}\frac{w(u)}{u^{1/4}}\mathop{\sum_{l,m,n}}_{nl\leq m\frac{u}{X}}\sum_{0<c\leq C_{0}^{\star}}\frac{S(nr_{1},-m,c)}{c}\left(\frac{c}{mn}\right)^{\frac{1}{4}}e\left(-2\sqrt{\frac{mnu}{c}}\right)w\left(\frac{mc}{n}\frac{u}{X}\right)\ll X^{\frac{3}{4}+\varepsilon}\left(\frac{X}{H}\right)^{\frac{3}{2}}.$ (8.6)
@@
$R_{3}(M,N,C,X,u)$ $=\mathop{\sum_{m}\sum_{n}}_{nl\leq m\frac{u}{X}}b_{1}\left(\frac{m}{M}\right)b_{2}\left(\frac{n}{N}\right)\sum_{c}\frac{S(nr_{1},-m,c)}{c}e\left(-2\sqrt{\frac{mnu}{c}}\right)w\left(\frac{mc}{n}\frac{u}{X}\right)V\left(\frac{c}{C}\right)g^{\star}(m,n,c),$
@@
We then have the following proposition which can be compared with Prop. 8.1.
@@
We give the details of the proof of Prop. 8.1 below and the proof of Prop. 8.2 is almost identical and we only give a very brief sketch of it at the end.
@@ 9.1. Decomposition of $R_{1}(M,N,C,X)$
say, where $\check{f}_{m,n,C}(\eta)$ is defined in (2.2). From the support of $V$, the argument $t$ of $f_{m,n,C}(t)$ (see (9.1)) is very small. Indeed, it lies in the the interval
@@
since $C>C_{0}$ (see (8.3)), $H\geq\sqrt{X}$ (see (3.3)) and $r=O(X^{1/3})$ (see Remark 15.1). We now expand $K_{2i\eta}(t)$ as a power series Reference?:
@@
$\mathcal{D}_{\pm}(M,N,C)=\mathop{\sum_{m}\sum_{n}}_{nl\leq m}b_{1}\left(\frac{m}{M}\right)b_{2}\left(\frac{n}{N}\right)\sum_{j=1}^{\infty}\rho_{j}(nr_{1})\rho_{j}(m)\check{f}_{m,n,C,\pm}(\kappa_{j}),$ (9.3)
$\mathcal{C}_{\pm}(M,N,C)$ $=\frac{1}{\pi}\mathop{\sum_{m}\sum_{n}}_{nl\leq m}b_{1}\left(\frac{m}{M}\right)b_{2}\left(\frac{n}{N}\right)\int\limits_{-\infty}^{\infty}(nmr_{1})^{-i\eta}\sigma_{2i\eta}(nr_{1})\sigma_{2i\eta}(m)\frac{\cosh(\pi\eta)\check{f}_{m,n,C,\pm}(\eta)}{\left|\zeta(1+2i\eta)\right|^{2}}\textnormal{d}\eta,$ (9.4)
@@
$\check{f}_{m,n,C,\pm}(\eta)$ $=\frac{4}{\pi}\int\limits_{0}^{\infty}K_{2i\eta,\pm}(t)f_{m,n,C}(t)\frac{\textnormal{d}t}{t}.$ (9.5)
@@
$I_{\pm}(m,n,C,\nu,\eta,r_{1})$ $=\int t^{\mp 2i\eta+2\nu}w\left(\frac{4\pi m\sqrt{mnr_{1}}}{nt}\right)V\left(\frac{4\pi\sqrt{mnr_{1}}}{tC}\right)g\left(m,n,\frac{4\pi\sqrt{mnr_{1}}}{t}\right)e\left(\frac{-X^{\frac{1}{2}}t^{\frac{1}{2}}(mn)^{\frac{1}{4}}}{\sqrt{\pi}r_{1}^{1/4}}\right)\frac{\textnormal{d}t}{t}.$
@@ 10. Estimation of $\mathcal{D}_{+}(M,N,C)$ and $\mathcal{C}_{+}(M,N,C)$
Here, the phase function $h_{+}$ has no stationary point and the first derivative bound together with the Weyl law suffices.
@@
[…]
@@ 10.1. Estimation of $\mathcal{D}_{+}(M,N,C)$
$\Theta_{c_{1},c_{2}}(M,N,K_{1},K_{2},\lambda)$ $=\sum_{\begin{subarray}{c}c_{1}K_{1}\leq\kappa_{j}\leq c_{2}K_{2}\end{subarray}}\mathop{\sum_{m}\sum_{n}}b_{1}\left(\frac{m}{M}\right)b_{2}\left(\frac{n}{N}\right)|\rho_{j}(nq)\rho_{j}(m)|\frac{e^{-\pi|\kappa_{j}|}}{\kappa_{j}^{\lambda}}$
[…]
@@
We subdivide the $j$-sum into dyadic pieces with $\kappa_{j}\asymp K$ such that $c_{1}K_{1}\leq K\leq c_{2}K_{2}$, and apply the Cauchy-Schwarz inequality to both the sums over $m$ and $n$, getting
@@
Since we know from Prop. 10.1 that the integral $I_{+}(m,n,C,\nu,\eta,r_{1})$ (see (9.8)) is negligibly small for $\kappa_{j}\gg\frac{MCX^{\varepsilon}}{NH}$, it is enough to estimate $\mathcal{D}_{+}(M,N,C)$ (see (9.3) and (9.7)) for $1\ll\kappa_{j}\ll\frac{MCX^{\varepsilon}}{NH}$. Using (9.3), Prop. 10.1, Prop. 10.3, and (8.4), we easily get
@@ 10.2. Estimation of $\mathcal{C}_{+}(M,N,C)$
$\Theta(M,N,K,\lambda)$ $=\mathop{\sum_{m}\sum_{n}}_{nl\leq m}b_{1}\left(\frac{m}{M}\right)b_{2}\left(\frac{n}{N}\right)\int\limits_{0}^{\infty}(nr_{1}m)^{-i\eta}\sigma_{2i\eta}(nr_{1})\sigma_{2i\eta}(m)\frac{e^{-\pi|\eta|}\cosh(\pi\eta)}{\eta^{\lambda}\left|\zeta(1+2i\eta)\right|^{2}}\textnormal{d}\eta$
@@
We know from [Tit86, (2.1.16)] that for $0<\eta<\varepsilon$ and also from [Tit86, (3.6.5)], for $\eta\geq\varepsilon$ respectively,
$\eta^{\lambda}\zeta(1+2i\eta)^{2}=\eta^{\lambda}\left(\frac{1}{2i\eta}+\gamma+O\left(\eta\right)\right)^{2}\gg\eta^{\lambda-2},\textnormal{ and }\frac{1}{\left|\zeta(1+2i\eta)\right|}\ll_{\varepsilon}|\log\eta|^{7}.$
@@
This is clear from Prop. 10.1, (9.4) and Prop. 10.5. ∎
@@ 11. Estimation of $\mathcal{D}_{-}(M,N,C)$ for $C_{0}\leq C\leq C_{1}$
Our next task is to estimate $\mathcal{D}_{-}(M,N,C)$ and $\mathcal{C}_{-}(M,N,C)$ (see (9.3), (9.4) and (9.6)). The sum $\mathcal{D}_{-}(M,N,C)$ (see (9.3)) is considerably more difficult to estimate than $\mathcal{D}_{+}(M,N,C)$ because the phase function $h_{-}(u)$ (see (9.10)) can have a stationary point inside the range of the integral $I_{-}(m,n,C,\nu,\eta,r_{1})$ (see (9.8)). Differentiating $h_{-}(u)$ (see (9.10)),
@@ 11.1. Further truncation of the $c$-sum
From (3.5) and the fact that $V$ is supported in $[1,4]$, it follows that $w\left(\frac{4\pi m\sqrt{mn}}{nu^{2}}\right)=1$ if $C_{0}<C\leq C_{1}$. In this range of $C$, the integral we need to analyze becomes
$I_{-}(m,n,C,\nu,\eta,r_{1})=2r_{1}^{\nu+i\eta}\int_{{\mathbb{R}}}A(u)e(-h_{-}(u))\textnormal{d}u,\textnormal{ where }A(u)=\frac{u^{4\nu}}{u}V\left(\frac{4\pi\sqrt{mn}}{u^{2}C}\right)g\left(m,n,\frac{4\pi\sqrt{mn}}{u^{2}}\right).$
@@
Unlike the previous case of $\mathcal{D}_{+}(M,N,C)$, an application of the Weyl law is not sufficient here. We need to obtain an explicit asymptotic expansion of the integral $I_{-}(m,n,C,\nu,\eta,r_{1})$ by applying Lemma 2.12. This is essential because if we estimate $I_{-}(m,n,C,\nu,\eta,r_{1})$ then our final bound will not be satisfactory. We will see that it is enough to consider the leading term $\mathcal{L}(m,n,\kappa_{j},\nu)$ of this asymptotic expansion (see (LABEL:lt). The phase $e(-h(u_{0}))$ occurring in $\mathcal{L}(m,n,\kappa_{j},\nu)$ simplifies to $(mn)^{-i\kappa_{j}}$, up to multiplication by a non-oscillatory function that does not involve $m$ and $n$, and we need to take advantage of the cancellations arising out of this oscillatory function. In order to implement this idea we will first have to separate the variables $m$ and $n$ which are at the moment entangled in the function $V$ and they are also constrained by the condition $nl\leq m$. Dropping the condition $nl\leq m$ by positivity as we had done earlier will be too costly here. We remove this condition by approximating the indicator function $\mathbf{1}_{[1,\infty)}(m/nl)$ by the Perron formula. Furthermore, the variables $m$ and $n$ occurring in the argument of the function $V$ are separated by Mellin inversion. Finally, after the two variables are separated, we apply the Cauchy-Schwarz inequality and the the Large Sieve inequality Lemma 2.6 for twisted Fourier coefficients $\rho_{j}(n)n^{-i\kappa_{j}}$.
@@
where $A$ and $\delta$ are arbitrary subject to the conditions $A>0$ and $0<\delta<1/10$. We define
@@
$\mathcal{D}_{-,k}(M,N,C)$ $=4\mathop{\sum_{m}\sum_{n}}_{nl\leq m}b_{1}\left(\frac{m}{M}\right)b_{2}\left(\frac{n}{N}\right)\sum_{j=1}^{\infty}\rho_{j}(nr_{1})\rho_{j}(m)\frac{1}{\sinh(2\pi\kappa_{j})\Gamma(1+2i\kappa_{j})}$ (11.8)
$\times\sum_{\nu=0}^{\infty}\frac{2^{-2\nu}}{\nu!(1+2i\kappa_{j})\dots(\nu+2i\kappa_{j})}\left[\frac{2r_{1}^{\nu+i\eta}}{\sqrt{h^{\prime\prime}_{-}(u_{0})}}e\left(-h_{-}(u_{0})-\frac{1}{8}\right)p_{k}(u_{0})\right],$
@@
$\mathcal{D}_{-,Z}(M,N,C)$ $=4\mathop{\sum_{m}\sum_{n}}_{nl\leq m}b_{1}\left(\frac{m}{M}\right)b_{2}\left(\frac{n}{N}\right)\sum_{j=1}^{\infty}\left|\frac{\rho_{j}(nr_{1})\rho_{j}(m)\kappa_{j}^{-A}}{\sinh(2\pi\kappa_{j})\Gamma(1+2i\kappa_{j})}\right|$ $\sum_{\nu=0}^{\infty}\frac{2^{-2\nu}r_{1}^{\nu}}{\nu!\left|(1+2i\kappa_{j})\dots(\nu+2i\kappa_{j})\right|}.$ (11.9)
@@
$\mathcal{L}(m,n,\kappa_{j},\nu,C)=2\frac{A(u_{0})}{\sqrt{h^{\prime\prime}_{-}(u_{0})}}e\left(-h_{-}(u_{0})-\frac{1}{8}\right)$ (11.10)
@@
$\frac{\pi}{2}\left(\frac{mnX}{C}\right)^{\frac{1}{2}}\leq\kappa_{j}\leq\pi\left(\frac{mnX}{C}\right)^{\frac{1}{2}};\quad\textnormal{in particular}\quad\frac{\pi}{2}N\left(\frac{Xl}{C}\right)^{\frac{1}{2}}\leq\kappa_{j}\leq 4\pi\left(\frac{MNX}{C}\right)^{\frac{1}{2}}.$ (11.11)
@@
$\mathcal{D}_{-,0}(M,N,C)=4\mathop{\sum_{m}\sum_{n}}_{nl\leq m}b_{1}\left(\frac{m}{M}\right)b_{2}\left(\frac{n}{N}\right)T(m,n,C),$ (11.12)
@@
$T(m,n,C)=\sum_{\nu=0}^{\infty}\sum_{\frac{\pi}{2}N\left(\frac{Xl}{C}\right)^{\frac{1}{2}}\leq\kappa_{j}\leq 4\pi\left(\frac{MNX}{C}\right)^{\frac{1}{2}}}\rho_{j}(nr_{1})\rho_{j}(m)\mathcal{V}(\kappa_{j},\nu,r_{1})\mathcal{L}(m,n,\kappa_{j},\nu,C),$
@@
$\mathcal{V}(\kappa_{j},\nu,r_{1})$ $=\frac{1}{\sinh(2\pi\kappa_{j})}\frac{2^{-2i\kappa_{j}-2\nu}r_{1}^{\nu+i\kappa_{j}}}{\Gamma(1+2i\kappa_{j})\nu!(1+2i\kappa_{j})\dots(\nu+2i\kappa_{j})}\ll\frac{e^{-\pi\kappa_{j}}r_{1}^{\nu}}{\nu!\sqrt{\kappa_{j}}},$ (11.13)
@@
$\tilde{T}(m,n,C)=\sum_{\nu=0}^{\infty}\sum_{\frac{\pi}{2}N\left(\frac{Xl}{C}\right)^{\frac{1}{2}}\leq\kappa_{j}\leq 4\pi\left(\frac{MNX}{C}\right)^{\frac{1}{2}}}|\rho_{j}(nr_{1})\rho_{j}(m)\mathcal{V}(\kappa_{j},\nu,r_{1})\mathcal{L}(m,n,\kappa_{j},\nu,C)|.$ (11.14)